# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ 3 Problem Setup
  • treatment $T(r)\in\mathcal{T}$, where $T(r)=1$ indicates that the treatment decision was $1$ when the recommendation was $r$; and
@@
Regarding fairness, we are concerned about disparities in utility and treatment benefits (resources or burdens) across different groups, denoted $A\in\{a,b\}$. (For notational brevity, we may generically discuss identification/estimation without additionally conditioning on the protected attribute.) For example, recommendations arise from binary high risk/low risk labels of classifiers. In practice, in consequential domains, classifier decisions are rarely automated but rather are used to inform humans-in-the-loop, who decide whether to assign treatment. For binary outcomes, we interpret $Y(t(r))=1$ as the positive outcome. When outcomes and treatments are binary, $Y\in\{0,1\},T\in\mathcal{T}$, where $\mathcal{T}=\{0,1\}$, we may further develop analogues of fair classification criteria. We let $c(r,t,y)\colon\{0,1\}^{3}\mapsto\mathbb{R}$ denote the cost function for $r\in\{0,1\},t\in\mathcal{T},y\in\{0,1\}$, which may sometimes be abbreviated $c_{rt}(y)$. We discuss identification and estimation on the basis of the following recommendation propensity $e_{r}$, treatment propensity $p_{t\mid r}$, and outcome $\mu_{t}$ models:
@@
We are generally instead interested in personalized recommendation rules, described via the policy function $\pi(r\mid X)\coloneqq\pi_{r}(X)$, which gives the probability of assignment of recommendation $r$ to covariates $X$. The average encouragement effect (AEE) is the difference in average outcomes if we refer everyone vs. no one, while the encouragement policy value $V(\pi)$ is the population expectation induced by the outcomes and treatment with recommendations following the policy distribution.
$AEE=\mathbb{E}[Y(T(1))-Y(T(0))],\qquad V(\pi)=\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$
We use the AEE terminology instead of ITT because the conventional first-stage intention-to-treat in ITT is actually our first-stage encouragement or recommendation. Because algorithmic decision-makers may be differentially responsive to recommendations and treatment effects may be heterogeneous, the optimal recommendation rule may differ from the (infeasible) optimal treatment rule when constraints are taken into account or for simpler policy classes.
Assumption 1 (Consistency and stable unit treatment values [SUTVA).
] $Y_{i}=Y_{i}(T_{i}(R_{i})).$
Assumption 2 (Conditional exclusion restriction).
@@
Assumption 4 (Stable responsivities under new recommendations).
$P(T=t\mid R=r,X)$ remains fixed from the observational to the future dataset.
Assumption 5 (Decomposable utilities).
$c(r,t,y)=c_{r}(r)+c_{t}(t)+c_{y}(y)$.
Our key assumption beyond standard causal inference assumptions is assumption˜2, the conditional exclusion restriction, i.e., that, conditional on observable information $X$, the recommendation has no causal effect on the outcome beyond its effect on the treatment probability. This assumes that all of the covariate information that is informative of downstream outcomes is measured. Although this assumption may not exactly hold in all applications, stating it is also a starting point for sensitivity analysis under violations of it (Kallus and Zhou, 2018; Kallus et al., 2019b).
Assumption˜4 is a structural assumption that limits our method to most appropriately reoptimize over small changes to existing algorithmic recommendations. For example, $p_{0\mid 1}(x)$ (disagreement with algorithmic recommendation) could be a baseline algorithmic aversion. Not all settings are appropriate for this assumption. We do not assume microfoundations on how or why human decision-makers deviate from algorithmic recommendations; rather, we take these patterns as given. Again, we can relax this assumption with sensitivity analysis. Assumption˜5 is a mild assumption on modeling utility, namely, that it is not defined on joint realizations of potential outcomes.
We first also assume overlap in recommendations and treatment. Later, however, we give robust methods for relaxing this assumption, leveraging our finer-grained characterization.
Assumption 6 (Overlap).
$\rho_{r}\leq e_{r}(X,A)\leq 1-\rho_{r};\;\;\rho_{t}\leq p_{t\mid r}(X,A)\leq 1-\rho_{t}\text{ and }\rho_{r},\rho_{t}>0$.
@@
Setting 1 (Randomized encouragement).
$R$ is (as-if) randomized and satisfies overlap (Assumption˜6).
@@
Then, $R$ can be interpreted as the ITT or prescription, whereas $T$ is the actual realization thereof. ˜1 models nonadherence situations where decision-makers can target encouragements but not direct receipt of the treatment itself.
Setting 2 (Algorithmic recommendation).
$R$ is the output of a predictive model and does not satisfy Assumption˜6.
We later extend our methods to the second setting, where $R$ does not satisfy overlap in recommendation but there is sufficient randomness in human decisions to satisfy overlap in treatment.
4 Analysis: What drives utility and budget unfairness under naive ITT targeting?
In order to motivate our later methodological work on estimating and optimizing encouragement policies under fairness constraints, we first answer the question: why is naive ITT targeting insufficient to understand the key drivers of unfairness? We focus on the class of threshold policies under global budget constraints. (Targeting encouragements is only needed when there is a budget constraint, as it is rare for encouragements to have negative treatment effects). We highlight that explicitly disaggregating effects on enrollment into treatment enables isolating the source of disparities into two substantively distinct mechanism: outreach vs. interventional effectiveness. We illustrate with a preview of our case study on SNAP outreach how unfairness surfaces in real-world scenarios. We develop analytical fairness decompositions that highlight key diagnostic metrics that lower-bound disparities in policy value. These metrics can be estimated from data — we highlight these in the case of the SNAP case study.
@@ 4 Analysis: What drives utility and budget unfairness under naive ITT targeting?
We begin by considering naive targeting on the intention-to-treat effect, i.e. viewing the encouragement as a treatment itself. Define the encouragement score:
$s(X,A)\coloneqq\tau(X,A)\,(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)).$
Proposition 1 (Quantile-threshold optimality).
@@
Assume the distribution of $s(X,A)$ is continuous (no atoms). Consider optimizing the ITT effect of encouragements under a global encouragement budget:
$\max_{\pi}\{\mathbb{E}[Y(\pi)]\colon\mathbb{E}[\pi(X,A)]\leq b\}$ (encouragement budget)
Let $q_{s}:[0,1]\to\mathbb{R}$ denote the (right–continuous) quantile function of $s$, $q_{s}(t)\;\coloneqq\;\inf\{z\in\mathbb{R}:\;\mathbb{P}\big(s(X,A)\leq z\big)\geq t\}$ Then the optimal policy is a threshold policy at the top-$b$ quantile of the encouragement score:
$\pi_{b}(x,a)\;=\;\mathbf{1}\!\left\{\,s(x,a)\geq\max(q_{s}(1-b),0)\,\right\}.$
Unfairness arises from differences in marginal levels, and joint dependence between treatment effect and nudgeability.
Fix a baseline policy $\pi_{0}$. The (group-conditional) improvement of $\pi$ over $\pi_{0}$ is
${Imp}_{a}(\pi;\pi_{0}):=\mathbb{E}\!\left[Y(\pi)-Y(\pi_{0})\mid A=a\right]=\mathbb{E}\!\left[\tau_{a}\,c_{a}\,(\pi(X,a)-\pi_{0}(X,a))\mid A=a\right]$
Improvement and disparity at a single threshold.
Choose $\pi_{0}=0$, i.e. encourage no-one. For example, this could be the status quo without any encouragements. Then
${Imp}_{a}(\pi^{*};0)=\mathbb{E}\!\left[s_{a}\,\mathbb{I}\{s_{a}\geq t^{*}\}\mid A=a\right],$
and analogously for $A=b$. Let the induced within-group budget be $\beta_{a}^{t^{*}}:=P(s_{a}\geq t^{*}\mid A=a)$.
The improvement at a threshold policy is driven by marginal levels of thresholded $\tau,c$ as well as their conditional covariance:
${Imp}_{a}(\pi^{*};0)={\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{*}]\ \mathbb{E}[c_{a}\mid s_{a}\geq t^{*}]}+\operatorname{Cov}\!\Big(\tau_{a}-\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{\star}],\ c_{a}-\mathbb{E}[c_{a}\mid s_{a}\geq t^{\star}]\,\Big|\ s_{a}\geq t^{\star}\Big)$
The conditional covariance term reflects the contribution of joint dependence of the treatment effect and compliance effects. It is zero (within the encouraged stratum) iff the treatment effect and compliance are uncorrelated, and positive when higher-benefit units also tend to comply more (as a “win-win" situation). Targeting encouragements is particularly beneficial when the people who benefit most are also the ones where encouragement nudges into treatment. Misalignment between effectiveness and nudgeability, surfacing as low or even negative dependence, suggest that encouragements need to be redesigned to achieve better targeting.
The disparity in improvement at $t^{*}$ is
$\Delta^{\mathrm{imp}}(t^{*}):={Imp}_{a}(t^{*})-{Imp}_{b}(t^{*})=\beta_{a}^{t^{*}}\mathbb{E}\!\left[s_{a}\,\mathbb{I}\{s_{a}\geq t^{*}\}\mid A=a\right]-\beta_{b}^{t^{*}}\mathbb{E}\!\left[s_{b}\,\mathbb{I}\{s_{b}\geq t^{*}\}\mid A=b\right]$
In the general case, the presence of disparity will depend on the budget (and therefore threshold): where in each group’s score distribution does the threshold fall, and what is the dependence structure of heterogeneity and nudgeability for each group. Differing strengths of dependence across groups can also result in disparities in group-specific utilities. The fundamental object that describes such dependence is the copula, the bivariate joint distribution between each group’s distribution of heterogeneous treatment effect and compliance effect functions. A general scale-free measure of dependence is the Spearman’s rank correlation, which measures the correlation of the ranks of two random variables and hence indicates whether a monotonic (but not necessarily linear) relationship exists between them. Under stronger structure on the copula, one can leverage disparities in metrics like Spearman’s rank correlation to conclude that disparities would persist uniformly over all potential thresholds. The required condition is a concordance-ordered copula family assumption, which posits that each group’s joint distribution of $(\tau,c)$ is indexed by a single dependence parameter $\theta$, such that larger $\theta$ values indicate stronger positive dependence—greater alignment between individuals who benefit most and those most likely to comply. Examples include the Gaussian or Archimedean (Gumbel, Clayton) copulas.
Theorem 3 (Spearman lower bound for disparity at the pooled cut).
Assume that for each group the copula belongs to a concordance-ordered family, indexed by a parameter $\theta,$ $\{C_{\theta}:\theta\in\Theta\}$, and that Spearman’s rank correlation $\rho_{\mathrm{Sp}}(\tau_{a},c_{a})$ (resp. $\rho_{\mathrm{Sp}}(\tau_{b},c_{b})$) is strictly increasing in $\theta$. Then there exist strictly positive constants $\kappa(t^{*})$, depending on the groupwise score distributions and copulas, so that
$\Delta^{\mathrm{imp}}(t^{*})\ \geq\ \kappa_{*}(t^{*})\ (\rho_{\mathrm{Sp}}^{\,a}-\rho_{\mathrm{Sp}}^{\,b})\ +(\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{*}]\mathbb{E}[c_{a}\mid s_{a}\geq t^{*}]-\mathbb{E}[\tau_{b}\mid s_{b}\geq t^{*}]\mathbb{E}[c_{b}\mid s_{b}\geq t^{*}])\ .$
In short: under concordance-ordered copulas, if one group has consistently stronger rank-alignment between treatment effect and compliance, then regardless of the budget threshold, disparities in policy value improvement persist. Such disparities could point to the need for more intensive outreach or differently designed encouragements.

### v5

@@ 3 Encouragement model and identification
  • treatment $T(r)\in\{0,1\}$, where $T(r)=1$ indicates that the treatment decision was $1$ when the recommendation was $r$; and
@@
We let $\pi$ denote a personalized encouragement recommendation rule. In general, for theoretical characterizations of optimal policies, we allow $\pi$ to depend on both covariates and protected attributes, and write
$\pi(x,a):=P(R=1\mid X=x,A=a),\qquad\pi_{r}(x,a):=P(R=r\mid X=x,A=a),$
When discussing practically implementable policy learning, however, we typically restrict attention to covariate-only policies $\pi(x)$. We seek to evaluate and optimize the value under different candidate policies $\pi$, which differ from the distribution of encouragements in our observed data.
Finally, for the purposes of assessing algorithmic fairness, we define the group-level protected attribute information $A\in\{a,b\}.$ The protected attribute $A$ designates group information by which it may be relevant to decompose groupwise performance disparities, such as race, gender, or age.
Our multi-objective criteria: treatment effectiveness and take-up.
The average encouragement effect (AEE) is the difference in average outcomes if we refer everyone vs. no one, while the encouragement policy value $V(\pi)$ is the population expectation induced by the outcomes and treatment with recommendations following the policy distribution.
$AEE\coloneqq\mathbb{E}[Y(T(1))-Y(T(0))],\qquad Y(\pi)\coloneqq\sum_{r\in\{0,1\}}\pi_{r}(X,A)Y(T(r)),\qquad V(\pi)\coloneqq\mathbb{E}[Y(\pi)]$
Regarding fairness, we are concerned about disparities in policy value and treatment take-up (resources or burdens) across different groups, denoted $A\in\{a,b\}$. (For notational brevity, we may generically discuss identification/estimation without additionally conditioning on the protected attribute.) Because the policy controls encouragement recommendations $R$ rather than treatment $T$ directly, fairness can be evaluated on several margins. Later, our primary focus is introducing awareness of disparities in access, take-up or demand for treatment.
$\Delta_{T}(\pi)\coloneqq\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]$ (1)
This is the natural access disparity in encouragement settings, because the planner controls recommendation rather than treatment directly. Other disparity measures that are useful additional diagnostics include: groupwise differences in final outcomes, treatment take-up, encouragement rates, and improvement from targeting — we discuss these in the context of our case studies later on.
In this paper, we first recommend auditing unconstrained policy optimization for disparities. Auditing can surface undesirable disparities, and we also consider multi-objective and budget-constrained policy optimization problem, where $B$ is a budget fraction of the population that can receive treatment:
Auditing: $\Delta_{T}(\pi^{*}),\text{ where }\pi^{*}\in\arg\max\mathbb{E}[V(\pi)]$
Fair policy optimization: $\max\left\{\mathbb{E}[Y(\pi)]-\lambda\left|\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\right|\colon\mathbb{E}[\pi(X)]\leq B\right\}$
The budget on encouragement is part of our default formulation since resource-constrained settings benefit the most from personalized targeting. Though the fairness-penalized and multi-objective version is practical, for theoretical analysis we often study the fairness-constrained problem. Since fairness is highly contextual, we generally don’t recommend trying to impose all disparity measures as constraints simultaneously. Instead, we use disparities as diagnostic measures to identify different gaps and to inform potentially different operational remedies, as illustrated in our later case studies.
Causal identification and estimation: key assumptions
We discuss identification and estimation on the basis of the following recommendation propensity $e_{r}$, treatment propensity $p_{t\mid r}$, and outcome model $\mu_{t}$:
@@
An important object for interpretation, appearing in the policy learning characterizations, is the effect of encouragement on treatment take-up.
$\kappa(x,a)\coloneqq p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)$
The following causal assumptions comprise our primary model, and our later extension to algorithmic advice is based on sensitivity analysis to some of them.
{assumption}
[Consistency and stable unit treatment values (SUTVA)] $Y_{i}=Y_{i}(T_{i}(R_{i})).$
{assumption}
[Exclusion / no direct effect of recommendation] $Y_{i}(t(r))=Y_{i}(t),\forall i,t,r$
{assumption}
@@
{assumption}
[Stable responsivities under new recommendations] $P(T=t\mid R=r,X)$ remains fixed from the observational to the future dataset. Our key assumption beyond standard causal inference assumptions is section 3, the conditional exclusion restriction/no direct effect assumption that the encouragement/recommendation has no causal effect on the outcome beyond its effect on the treatment. This assumption is typically mild for encouragements, although it requires that all of the covariate information that is informative of downstream outcomes is measured.
We begin with Section 3 first to highlight the two distinct mechanisms at play, compliance and treatment response, and discuss potential violations later on.
Section 3 is a structural assumption that limits our method to most appropriately reoptimize over small changes to existing algorithmic recommendations. For example, $p_{0\mid 1}(x)$ (disagreement with algorithmic recommendation) could be a baseline algorithmic aversion. Not all settings are appropriate for this assumption. We do not assume microfoundations on how or why human decision-makers deviate from algorithmic recommendations; rather, we take these patterns as given.
Under these assumptions, we obtain causal identification of the policy value via outcome modeling (sometimes called regression adjustment). Causal identification rewrites the causal estimand in terms of probability distributions estimable from data. The argument follows by applying the conditional exclusion restriction and consistency but, crucially, does not rely on overlap.
Proposition 3.1 (Regression adjustment identification)
$\mathbb{E}[Y(\pi)]$ $\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)p_{t\mid r}(X)].$
We first also assume overlap in recommendations and treatment. {assumption}[Overlap] $\rho_{r}\leq e_{r}(X,A)\leq 1-\rho_{r};\;\;\rho_{t}\leq p_{t\mid r}(X,A)\leq 1-\rho_{t}\text{ and }\rho_{r},\rho_{t}>0$.
@@
Setting 3.2 (Setting 1: Randomized encouragement)
The encouragement $R$ is (as-if) randomized and satisfies overlap (Proposition 3.1).
@@
Then, $R$ can be interpreted as the ITT or prescription, whereas $T$ is the actual realization thereof. 3.2 models nonadherence situations where decision-makers can target encouragements but not direct receipt of the treatment itself.
Our perspective: Fairness constraints in causal intention-to-treat (ITT) analyses.
We consider covariate-conditional intention-to-treat causal effect under the assumption of treatment unconfoundedness. We evaluate encouragement policies by their population-level value and by groupwise impacts under the treatment behavior they induce. Thus, treatment take-up is central for fairness constraints, while utility is assessed through policy value; we do not pursue IV or principal-stratum estimands.
3.1 Policy characterizations
@@
Next, we focus on basic characterizations of the optimal encouragement policy under a constrained resource budget, where targeting becomes particularly useful. We begin by considering naive targeting on the intention-to-treat effect, i.e. viewing the encouragement as a treatment itself. The optimal policy thresholds on what we call the encouragement score:
$s(X,A)$ $\coloneqq\mathbb{E}[Y\mid R=1,X,A]-\mathbb{E}[Y\mid R=0,X,A]$
$=\mathbb{E}[Y(T(1))-Y(T(0))\mid X,A].$ (under sections 3 and 3)
$=\mathbb{E}[(Y(1)-Y(0))(T(1)-T(0))\mid X,A].$ ( section 3)
Define a generic measure of treatment efficacy $\theta(X,A)$, akin to a conditional Wald estimand, that can be used to define a generic decomposition of the encouragement score as follows:
$\theta(X,A)\coloneqq\frac{s(X,A)}{\mathbb{E}[T(1)-T(0)\mid X,A]},\;\;s(X,A)\coloneqq\theta(X,A)\,(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)).$
We define the heterogeneous treatment effect $\tau(X,A)=\mathbb{E}[Y(1)-Y(0)\mid X,A].$ Crucially, Section 3 enables interpretation of the treatment efficacy part of the decomposition of $s(X,A)$ as a heterogeneous treatment effect, i.e. $\theta(X,A)=\tau(X,A)$, so that
$s(X,A)=\mathbb{E}[Y(1)-Y(0)\mid X,A]\,(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))=\tau(X,A)\kappa(X,A)$ (Section 3, Section 3)
We study decompositions of the encouragement score to interpret optimal policies, where naive ITT targeting by default optimizes it.
Proposition 3.3 (Quantile-threshold optimality)
@@
Assume the distribution of $s(X,A)$ is continuous (no atoms). Consider optimizing the ITT effect of encouragements under a global encouragement budget: $\pi^{*}_{\beta}(x,a)\in\arg\max_{\pi}\{\mathbb{E}[Y(\pi)]\colon\mathbb{E}[\pi(X,A)]\leq\beta\}$ Let $t_{\beta}\coloneqq\max\!\left\{\inf\{z:P(s(X,A)\leq z)\geq 1-\beta\},\,0\right\}$. Then the optimal policy is a threshold policy at the top-$\beta$ quantile of the encouragement score (so long as it is positive):
$\pi^{*}_{\beta}(x,a)\;=\;\mathbf{1}\!\left\{\,s(x,a)\geq t_{\beta}\,\right\}$ (2)
Threshold policies on encouragement score solve budget-constrained naive targeting; later we study properties of such threshold policies.
3.2 Beyond the base assumptions
Beyond recommendation overlap (Proposition 3.1)
Although we begin with these base assumptions, we can handle violations of certain of these assumptions with separate robustness checks. The most central of these robustness analyses is our extension to algorithmic advice by allowing for violations of Proposition 3.1, for which we discuss robust optimization
Setting 3.4 (Setting 2: Algorithmic recommendation)
The algorithmic recommendation $R$ is the output of a predictive model and does not satisfy Proposition 3.1.
We later extend our methods to the second setting, where $R$ does not satisfy overlap in recommendation but there is sufficient randomness in human decisions to satisfy overlap in treatment. For example, such recommendations arise from binary high risk/low risk labels of classifiers since decisions in consequential domains are rarely automated but rather are used to inform humans-in-the-loop, who decide whether to assign treatment.
Beyond treatment unconfoundedness (Section 3)
It may look strange to reference the heterogeneous treatment effect in a paper on encouragement designs/adherence. The key is our Section 3, on unconfounded treatments. Other papers making similar assumptions include Rahier et al. (2021) in e-commerce, Imai et al. (2010) in mediation analysis, and Rudolph et al. (2020).
Framework Score decomposition ($s(x,a)=$) Interpretation of multiplier
Unconfounded treatments (A4) $\kappa(x,a)\tau(x,a)$ Under Section 3, $\tau(x,a)=\mathbb{E}[Y(1)-Y(0)\mid X=x,A=a]$, the conditional treatment effect.
IV + monotonicity $\kappa(x,a)\tau_{c}(x,a)$ $\tau_{c}(x,a)=\mathbb{E}[Y(1)-Y(0)\mid T(1)>T(0),X=x,A=a]$, the conditional LATE / CLATE.
Wang–Tchetgen Tchetgen $\kappa(x,a)\theta(x,a)$ $\delta_{W}(x,a)=\Delta_{Y}(x,a)/\kappa(x,a)$, the conditional Wald estimand; under A5.a or A5.b it coincides with the conditional ATE.
Table 1: Alternative interpretations of $\theta(x,a)$ in the encouragement score $s(x,a)$. Full assumptions are given in the appendix.
Section 3 may be violated due to the role of private information governing how humans make decisions. If one objects to our treatment unconfoundedness assumption, there are a few options. An alternative route available when the recommendation is truly randomized is to interpret the recommendation $R$ as an instrumental variable for treatment $T$. In that case, a similar decomposition holds but each of the treatment-take-up vs. treatment-effectiveness objects has a slightly different interpretation, outlined in Table 1. See Section 8.2 in the Appendix for more details.
In our Setting 2 (Algorithmic advice), such alternative identification approaches are not available. However, our paper focuses on data- and covariate-rich settings. It may be plausible that observed covariates explain most, but not all of selection into treatment - so that sensitivity analysis approaches (Kallus and Zhou 2021b, Yadlowsky et al. 2018) can robustly optimize over “last-mile” unobserved confounders. We illustrate this approach in our algorithmic advice case study on supervised release in Section 6.2 and how to conduct such a sensitivity analysis in the Appendix.
4 Audit: are disparities driven by take-up or efficacy?
In encouragement designs, value is generated through two distinct causal channels: encouragement changes treatment uptake, and treatment uptake changes outcomes. As a result, a policy that targets on the individual encouragement effect can generate group disparities for two substantively different reasons. First, groups may differ in treatment benefit. Second, groups may differ in responsiveness to encouragement. These mechanisms have different normative interpretations and different operational remedies, so we analyze them separately.
We proceed in three steps. First, we discuss how the relationship between treatment benefit and compliance is central to quantifying the improvements of personalized targeting, and therefore why separately analyzing treatment take-up and efficacy is crucial to analyze mechanisms underlying disparities in downstream improvements. Second, we show that even when treatment effects are distributionally identical across groups, naive ITT targeting can still generate persistent disparities through compliance-only covariates. Finally, we illustrate the value of auditing both take-up and efficacy separately in a real-world case-study where we highlight that such phenomena do appear empirically.
4.1 The joint dependence of treatment effects and compliance governs policy value improvements from targeting.
To see why it is important to distinguish between treatment efficacy and treatment take-up, consider a healthcare setting where patients are prescribed an overall beneficial medication, whose effects may nonetheless vary, but is potentially burdensome to obtain and administer. Suppose that a personalized prescription policy yields lower encouragement value for a particular group. This disparity could arise because the treatment is less effective for that group, or because patients are less likely to adhere to the prescription when encouraged. These two mechanisms imply fundamentally different responses: differences in treatment efficacy call for revisiting the underlying science of the treatment, whereas differences in take-up point to operational interventions such as adherence support, outreach design, or access improvements.
Targeting is most effective when individuals who benefit most are also those most responsive to encouragement. But there can be groupwise heterogeneity in both. We breakdown the groupwise-improvement from personalized ITT targeting to highlight different sources of disparities. Choose $\pi_{0}=0$, i.e. encourage no-one. For example, this could be the status quo without any encouragements. Then, for a group $A=a$, we can define the improvement and a decomposition into marginal effects and the conditional covariance between compliance and treatment efficacy scores:
${Imp}_{a}(\pi^{*};0)\coloneqq\mathbb{E}[Y(\pi)-Y(0)\mid A=a]=\mathbb{E}\!\left[s_{a}\,\mathbb{I}\{s_{a}\geq t^{*}\}\mid A=a\right]$
$=P\left(s_{a}\geq t^{*}\mid A=a\right)\left(E\left[\tau_{a}\mid s_{a}\geq t^{*},A=a\right]E\left[\kappa_{a}\mid s_{a}\geq t^{*},A=a\right]+\operatorname{Cov}\left(\tau_{a},\kappa_{a}\mid s_{a}\geq t^{*},A=a\right)\right)$
The conditional covariance term reflects the contribution of the joint dependence between treatment effects and compliance. It is zero within the encouraged stratum if treatment benefit and compliance are uncorrelated, and positive when individuals who benefit more from treatment also tend to be more responsive to encouragement. Thus, targeting is especially valuable when treatment efficacy and responsiveness are aligned. More generally, groups can differ in the gains they derive from personalized targeting because of differences either in the marginal levels of treatment benefit and compliance or in the alignment between them.
A stylized access-driven model of encouragement disparities.
Observed disparities in encouragement effects or improvements from personalized targeting can arise under very different mechanisms. We give a possibility result of low value-fairness trade-off arising due to naive ITT targeting exploiting “compliance proxies”, covariates that affect compliance alone but not treatment effects.
To concretize, consider a simple model with covariates partitioned as $(X^{s},X^{\kappa},X^{\tau})$ where $X$ is a shared factor affecting both compliance and heterogeneous treatment effect, $X^{\kappa}$ modifies compliance alone, and $X^{\tau}$ modifies the heterogeneous treatment effect alone. The protected attribute $A$ has no direct effect on compliance or the heterogeneous treatment effect, but $A$ impacts the distribution of a compliance-relevant covariate $X^{\kappa}$, i.e. $X^{\kappa}\mid A$ differs.
To interpret this example, $X^{\kappa}$ could be zip code that affects location and therefore distance to medical services and adherence. With other health information about a person, zip code wouldn’t plausibly biologically affect treatment effectiveness, but could affect adherence/compliance. Such compliance differences alone can generate persistent disparities in encouragement and treatment take-up. Thus, naive targeting may reward ease of nudging rather than treatment benefit, especially when it exploits covariates that proxy for access or adherence rather than efficacy. The resulting disparities can be normatively concerning when lower responsiveness reflects structural access barriers rather than lower need. Naive ITT targeting favors individuals who are easier to reach, not necessarily those for whom treatment is most valuable, and therefore could entrench structural disparities.
In Section 9.1 we analytically characterize the favorable fairness-value tradeoff in this instance. A simple proxy-blind targeting approach, i.e. optimizing over rules that do not use compliance proxy information, nearly eliminates group-wise encouragement disparity from $0.1232$ to $0.0088$ with a modest value loss of $0.0031$ (1.4%).
This stylized example shows that disparities under naive ITT targeting can arise entirely from differential compliance, even when treatment benefits are identical across groups, and therefore that fairness adjustments could nearly eliminate such disparities at potentially low tradeoff in overall value. Next, we apply such diagnostics on a real encouragement setting, where we see similar phenomena show up in real-world data.